OpenStudy (anonymous):

determine if diverge or coverge integral 3 to -3 y/sqrt 9-y^2

OpenStudy (anonymous):

\[ \int\limits_{-3}^{3} y/9-y^2\]

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

?

OpenStudy (anonymous):

yah

OpenStudy (anonymous):

no srry

OpenStudy (anonymous):

there is a square root sih around the 9- y^2

OpenStudy (anonymous):

Okay let see

OpenStudy (bahrom7893):

0

OpenStudy (anonymous):

u=9-y^2 du=-2y dy

OpenStudy (bahrom7893):

lol i dunno the rest, i can solve the integral, i suck at series, sorry...=(

OpenStudy (anonymous):

\[\sqrt(u)du\]

OpenStudy (anonymous):

can i also use u substitution

OpenStudy (anonymous):

\[-2\int\limits(du/\sqrt(u)\]

OpenStudy (anonymous):

i mean trig subsitution

OpenStudy (anonymous):

yes that's what I did

OpenStudy (anonymous):

Trig is more work

OpenStudy (anonymous):

yah i had a hard time going that route

OpenStudy (anonymous):

just use u-sub don't use trig sub unless you really have to

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

im kinda stuck on the last step when plug back the points